solve by using elimination method. x+4y=32
-x+8y=52

Answer by stanbon(75887)

solve by using elimination method.
x+4y=32
-x+8y=52
------------
Add and solve for "y":
12y = 84
y = 7
---------
solve for "x":
x + 4*7 = 32
x = 4
----------------------
Cheers,
Stan H.
===========